OpenStudy (anonymous):

can anyone explain to me the basic rules of function transformations/translation

OpenStudy (anonymous):

Subtract from the variable in question the distance you want to move. Subtract from x to move left or right, and from y to move up and down. Example: y = 2x + 3. To move it up 3 units, subtract 3 from y: y - 3 = 2x +3. Simplify to y = 2x + 6. To move it left 5 units, subtract -5 from x: y = 2(x+5) +3. Simplify to y = 2x + 13.